OpenStudy (anonymous):

Your annual salary was $26,300 in2002 and $29,700 in2004. Assume your salary can be modeled by a linear equation. (a) Write a linear equation giving your salary S in terms of the year. Let t : 2 represent 2002. (b) Use the linear model to predict your salary in 2008.

OpenStudy (anonymous):

In 2 years it increased by $3,400, which is an average of $1,700 per year. Assuming t =0 represents 2000, it would have been $26,300 - (2*1700) = $22,900 in 2002. So the linear model is S = 22900 + 1700t. In 2008, t = 8, so your salary should be 22900 + 1700*8 = ...

OpenStudy (anonymous):

sorry I don't know is it for (a) or (b)??

OpenStudy (anonymous):

(a) S = 22900 + 1700t (b) S = 22900 + 1700*8 = 22900 + 13600 = $36,500.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No worries. A medal would be nice though :-P